# Exploring the Potential of Large Language Models for Heterophilic Graphs

### v2

@@ 4.2 Performance Comparison
Table 2: Accuracy for node classification of different methods. (Best results are bolded; runners-up are underlined.)
@@
For different strategies of LLM4HeG, directly fine-tuning the SLMs often leads to notable performance decline compared to fine-tuning the LLM. However, the distilled SLMs attain performance comparable to that of the fine-tuned LLM, demonstrating the effectiveness of the distillation process in retaining heterophily-specific knowledge. The only exception is the performance on the Amazon dataset, where the directly fine-tuned SLMs achieves similar performance as the LLM and the distilled ones. This may be due to the semi-structured patterns within the textual descriptions in the Amazon dataset such as product specifications, which are relatively easy to capture by both LLMs and SLMs.
@@ 4.3 Model Analysis
Ablation Study.
To evaluate the effectiveness of the learnable weight for adaptive message passing and deep node features from LLM, we conducted experiments using different variants of our model: (1) w/o reweight: This variant only uses the graph-based weight $w_{uv}^{G}$ (2) w/o learnable weights: This variant uses fixed weights instead of learnable ones. Specifically, we manually set the weights for homophilic and heterophilic pairs to 1 and -1, respectively, as these values were found to perform well across most datasets.
@@
Analysis of edge discrimination by LLM/SLMs.
As the edge reweighting in Stage 2 depends on the effectiveness of edge discrimination in Stage 1, we further evaluate the edge classification performance (F1 score) of different models on the node pairs used for Stage 2. As shown in Table 3, compared with the fine-tuned LLM, directly fine-tuning the SLMs generally gives worse edge discrimination performance, which is reasonable considering the model capacity. However, the distilled SLM manages to maintain a comparable performance with only a marginal drop. This indicates the effectiveness of the distillation process, which allows the distilled SLM to retain heterophily-specific knowledge of the fine-tuned LLM. The results are consistent with the node classification performance reported in Table 2.
Efficiency study.
Fig. 3 illustrates the efficiency of our framework using an LLM or distilled SLMs in terms of training and inference times, measured in minutes. For the LLM approach, the training time denotes the fine-tuning time in Stage 1 on the selected node pairs (see Section 3.3). For the distilled SLMs, as discussed in Section 3.5, we employ the fine-tuned LLM to generate pseudo-labels for further fine-tuning the SLM. Thus, its training time includes fine-tuning the LLM, generating the pseudo-labels, and fine-tuning the SLM. Thus, the total training time for model distillation is slightly higher than the LLM. On the other hand, inference time includes predicting edge types using either the LLM or distilled SLM, as required by the edge reweighting module in Stage 2. It is worth noting that the inference times of SLMs are significantly lower than LLMs, especially for larger datasets such as Amazon. Hence, the distilled SLMs can be more easily deployed given their smaller size and faster inference time, while maintaining competitive performance as shown in Table 2.
4.3.1 Plug-and-play with various backbones.
Our method is designed to be highly flexible, allowing it to be integrated with various GNN backbones. To demonstrate the flexibility, we tested our approach with several well-known GNN architectures, including GCN, GAT, H2GCN and FAGCN. We provide the implementation details of the integrating process in Appendix B. As shown in Table 4, our method consistently enhances the performance of these backbones. This plug-and-play capability highlights the versatility of our approach, making it valuable for improving multiple backbones without the need for significant modifications.
@@ 5 Conclusion
Appendix
@@ Appendix A More Details for Datasets
  • Actor Pei et al. (2020); Tang et al. (2009) is an actor-only induced subgraph of the film-director-actor-writer network. In this graph, nodes represent actors, and an edge between two nodes indicates their co-occurrence on the same Wikipedia page. The classification task involves categorizing actors into five distinct classes based on their roles. We selected the actors based on category information provided in the metadata, focusing on those with high occurrence frequencies. The category keywords of the selected actors include “American film actors", “American film and television actors", “American stage and television actors", “English" and “Canadian". Afterward, we construct the graph based on the edges and remove the isolated nodes from the graph.
  • Amazon Platonov et al. (2023) is constructed from the Amazon product co-purchasing network metadata. In this dataset, nodes represent products such as books, music CDs, DVDs, and VHS video tapes. Edges link products that are frequently bought together. The goal here is to predict the average rating a product receives from reviewers, with ratings grouped into five classes. To manage the graph’s complexity, only the largest connected component of the 5-core of the graph is considered.
@@ Appendix B More Implementation Details
Node Pairs
In Stage 1, we train the edge discriminator by sampling node pairs from the graph, with the selection process tailored to the characteristics of each dataset. For small graphs like Cornell, Texas, and Wisconsin, we select all node pairs within the training set, including those without direct edges. For the Actor dataset, node pairs are selected based on the 1-hop and 2-hop neighbor relationship, while for the Amazon dataset, we focus on node pairs with a 1-hop neighbor relationship, taking into account the graph’s size.
In the distillation stage, we use node pairs from the validation and testing data as additional samples, allowing the fine-tuned LLM to generate pseudo-labels. For small graphs like Cornell, Texas, and Wisconsin, we select node pairs with 1-hop and 2-hop neighbor relationships, while for the Actor and Amazon datasets, we choose node pairs with 1-hop neighbor relationships. The number of node pairs used for training in Stage 1 and distillation is shown in Table 5.
@@
Table 5: The number of node pairs in Stage 1 and distillation.
Backbones.
We provide the implementation details for integrating the LLM4HeG with other backbones. As discussed in Sec. 3.4, the edge reweighting in Stage 2 combines the LLM-based weight with graph-based information. For the backbones that don’t contain an additional edge weight learning module (e.g., GCN Kipf and Welling (2016) and H2GCN Zhu et al. (2020)), we only use the edge weight obtained from LLM. For the backbones with specific designs of the edge weight (e.g., GAT Veličković et al. (2018)), we follow the Eq. (5) to combine the LLM-based weight and graph-based weight to perform fine-grained context aggregation of GNN.
@@ Appendix D More Details for Baselines
H2GCN Zhu et al. (2020) considers higher-order neighbors, ego-neighbor embedding separation and intermediate layer representations for heterophily graph.

### v3

@@ 4.2 Performance Comparison
Table 2: Accuracy for node classification of different methods. (Best results bolded; runners-up underlined.)
@@
For different strategies of LLM4HeG, directly fine-tuning the SLMs often leads to notable performance decline compared to fine-tuning the LLM. However, the distilled SLMs attain performance comparable to that of the fine-tuned LLM, demonstrating the effectiveness of the distillation process in retaining heterophily-specific knowledge. The only exception is the performance on the Amazon dataset, where the directly fine-tuned SLMs achieve similar performance as the LLM and the distilled ones. This may be due to the semi-structured patterns within the textual descriptions in the Amazon dataset such as product specifications, which are relatively easy to capture by both LLMs and SLMs. We also present experiments that directly utilize LLMs for node classification in Appendix E.
@@ 4.3 Model Analysis
Ablation Study. To evaluate the effectiveness of the learnable weight for adaptive message passing and deep node features from LLM, we conducted experiments using different variants of our model: (1) w/o reweight: This variant only uses the graph-based weight $w_{uv}^{G}$ (2) w/o learnable weights: This variant uses fixed weights instead of learnable ones. Specifically, we manually set the weights for homophilic and heterophilic pairs to 1 and -1, respectively, as these values were found to perform well across most datasets.
@@
Analysis of edge discrimination by LLM/SLMs. As the edge reweighting in Stage 2 depends on the effectiveness of edge discrimination in Stage 1, we further evaluate the edge classification performance (F1 score) of different models on the node pairs used for Stage 2. As shown in Table 3, compared with the fine-tuned LLM, directly fine-tuning the SLMs generally gives worse edge discrimination performance, which is reasonable considering the model capacity. However, the distilled SLM manages to maintain a comparable performance with only a marginal drop. This indicates the effectiveness of the distillation process, which allows the distilled SLM to retain heterophily-specific knowledge of the fine-tuned LLM. The results are consistent with the node classification performance reported in Table 2.
Performance on inductive setting. Our method is inherently adaptable to inductive node classification, where test nodes remain entirely unseen during training. We conduct additional experiments on inductive test sets, requiring the model to classify them based on their raw textual attributes and structural connections. As shown in Fig. 3, our approach consistently outperforms baseline methods in this setting. These results indicate that our method can effectively benefit from the semantic understanding of LLMs in both transductive and inductive settings.
Efficiency study. Fig. 4 illustrates the efficiency of our framework using an LLM or distilled SLMs in terms of training and inference times, measured in minutes. For the LLM approach, the training time denotes the fine-tuning time in Stage 1 on the selected node pairs (see Section 3.3). For the distilled SLMs, as discussed in Section 3.5, we employ the fine-tuned LLM to generate pseudo-labels for further fine-tuning the SLM. Thus, its training time includes fine-tuning the LLM, generating the pseudo-labels, and fine-tuning the SLM. Thus, the total training time for model distillation is slightly higher than the LLM. On the other hand, inference time includes predicting edge types using either the LLM or distilled SLM, as required by the edge reweighting module in Stage 2. It is worth noting that the inference times of SLMs are significantly lower than LLMs, especially for larger datasets such as Amazon. Hence, the distilled SLMs can be more easily deployed given their smaller size and faster inference time, while maintaining competitive performance as shown in Table 2.
4.3.1 Plug-and-play with various backbones
Our method is designed to be highly flexible, allowing it to be integrated with various GNN backbones. To demonstrate the flexibility, we tested our approach with several well-known GNN architectures, including GCN, GAT, H2GCN, FAGCN and GBK-GNN. We provide the implementation details of the integration process in Appendix B. Table 4 indicates that our method consistently enhances the performance of these backbones. This capability highlights the versatility of our approach, making it valuable for improving multiple backbones without the need for significant modifications.
@@ 4.3.1 Plug-and-play with various backbones
We also conduct experiments on the influence of the initial values of $w_{\text{Ho}}$ and $w_{\text{He}}$ in Eq. (4). The results in Table 5 show that the classification performance is relatively stable across different initial values of $w_{\text{Ho}}$ and $w_{\text{He}}$, with minor variations observed for specific datasets. This suggests that the model is robust to different initializations, and can generally find the optimal values for $w_{\text{Ho}}$ and $w_{\text{He}}$ despite different initializations.
$w_{\text{Ho}}$ $w_{\text{He}}$ Cornell Texas Wisconsin Actor Amazon
0.5 0.5 78.57±2.6 89.55±2.9 85.44±2.5 76.83±0.6 51.47±0.4
1 0 77.86±2.4 89.55±2.5 85.44±2.1 77.00±0.4 51.42±0.6
1.5 -0.5 77.62±2.9 89.09±3.3 86.14±2.1 76.82±0.5 51.53±0.4
2 -1 78.57±1.8 87.73±3.6 85.44±1.8 76.88±0.5 51.40±0.4
2.5 -1.5 78.57±1.5 87.95±2.3 86.14±1.8 76.99±0.5 51.30±0.3
Table 5: Effect of the initial weights of $w_{\text{Ho}}$ and $w_{\text{He}}$.
@@ 5 Conclusion
Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendices
@@ Appendix A More Details for Datasets
  • Actor Pei et al. (2020); Tang et al. (2009) is an actor-only induced subgraph of the film-director-actor-writer network. In this graph, nodes represent actors, and an edge between two nodes indicates their co-occurrence on the same Wikipedia page. The task involves categorizing actors into five distinct classes based on their roles. We selected the actors based on category information provided in the metadata, focusing on those with high occurrence frequencies. The category keywords of the selected actors include “American film actors", “American film and television actors", “American stage and television actors", “English" and “Canadian". Afterward, we construct the graph based on the edges and remove the isolated nodes from the graph.
  • Amazon Platonov et al. (2023) is constructed from the Amazon product co-purchasing network metadata. In this dataset, nodes represent products such as books, music CDs, DVDs, and VHS video tapes. Edges link products that are frequently bought together. The goal is to predict the average rating a product receives from reviewers, with ratings grouped into five classes. To manage the graph’s complexity, only the largest connected component of the 5-core of the graph is considered.
@@ Appendix B More Implementation Details
Node Pairs. In Stage 1, we train the edge discriminator by sampling node pairs from the graph, with the selection process tailored to the characteristics of each dataset. For small graphs like Cornell, Texas, and Wisconsin, we select all node pairs within the training set, including those without direct edges. For the Actor dataset, node pairs are selected based on the 1-hop and 2-hop neighbor relationship, while for the Amazon dataset, we focus on node pairs with a 1-hop neighbor relationship, taking into account the graph’s size.
In the distillation process, we use node pairs from the validation and testing data as additional samples, allowing the fine-tuned LLM to generate pseudo-labels. For small graphs like Cornell, Texas, and Wisconsin, we select node pairs with 1-hop and 2-hop neighbor relationships, while for the Actor and Amazon datasets, we choose node pairs with 1-hop neighbor relationships. The number of node pairs used for training in Stage 1 and distillation is shown in Table 6.
@@
Table 6: The number of node pairs in Stage 1 and the distillation process.
Backbones. We provide the implementation details for integrating the LLM4HeG with other backbones. As discussed in Sec. 3.4, the edge reweighting in Stage 2 combines the LLM-based weight with graph-based information. For the backbones that don’t contain an additional edge weight learning module (e.g., GCN Kipf and Welling (2016) and H2GCN Zhu et al. (2020)), we only use the edge weight obtained from LLM. For the backbones with specific designs of the edge weight (e.g., GAT Veličković et al. (2018), FAGCN Bo et al. (2021) and GBK-GNN Du et al. (2022)), we follow the Eq. (6) to combine the LLM-based weight and graph-based weight to perform fine-grained context aggregation of GNN.
@@ Appendix D More Details for Baselines
  • H2GCN Zhu et al. (2020) considers higher-order neighbors, ego-neighbor embedding separation and intermediate layer representations for heterophilic graph.
@@
  • GBK-GNN Du et al. (2022) introduces a learnable kernel selection gate to discriminate node pairs and apply two different kernels for homophily and heterophily node pairs.
@@
Appendix E Comparison to Direct LLM Predictions
Given the rich textual data associated with the nodes, it is possible to directly feed this information into the LLM for classification, bypassing the use of a GNN. We conduct experiments using the original and fine-tuned LLMs.
(1) “LLM-nofinetune” employs the original Vicuna 7B model to make the prediction, with the following prompt template for the Cornell, Texas, and Wisconsin datasets. A similar prompt template with different background and task descriptions is used for the Actor and Amazon datasets.
Background: I have a dataset containing web page information collected from computer science department websites of various universities. These web pages have been manually categorized into five categories, including student, staff, faculty, course, and project. Task: I will provide you with the text information of a web page, and I would like you to classify it into one of the following categories: student, staff, course, faculty, or project. The web page content: <text> You may only output the category name, and do not discuss anything else!
(2) “LLM-finetune” employs the fine-tuned Vicuna 7B model using the edge discriminator in Stage 1 to make the prediction, using the same prompt template as above.
As shown in Table 7, the performance of the original LLM without fine-tuning performs poorly when directly provided with textual information. This is likely because while LLMs are powerful for natural language processing tasks, they struggle to infer node categories without adaptation to the specific task. After fine-tuning with the LLM-enhanced edge discriminator in Stage 1, the performance improves significantly, but is still worse than LLM4HeG. This highlights the importance of combining node semantic features, structural contexts, and LLM-inferred edge characteristics for effective node classification.
Cornell Texas Wisconsin Actor Amazon
LLM-nofinetune 26.16 25.00 26.32 24.75 35.44
LLM-finetune 61.90 40.91 71.93 59.96 36.52
LLM4HeG 75.95 87.50 84.91 76.54 51.53
Table 7: Performance comparison with LLMs.